OpenStudy (anonymous):

Use the elimination method to find the solutions to the system. Write solutions as points (x,y) x^2+y^2=4 2x^2-y^2=8 (x,y)= ( ) and ( )

myininaya (myininaya):

3x^2=12 x^2=4 x=2,-2 let me if you need more help

OpenStudy (anonymous):

no that is what I had come up with as well thank you

OpenStudy (anonymous):

both ways y=0 correct?

myininaya (myininaya):

so you have x^2+y^2=4 y^2=4-x^2 y^2=4-4=0 y=0 yep (2,0) and (-2,0)

OpenStudy (anonymous):

awesome thank you so much

myininaya (myininaya):

np